# Electricity Demand and Grid Impacts of AI Data Centers: Challenges and Prospects

### v3

@@ III-B AI Computing Load Patterns at Different Stages
Since AI computing constitutes the dominant share of electricity consumption in AI data centers, this subsection analyzes its load patterns and characteristics across the four main stages of AI model workflows: preparation, training, fine-tuning, and inference [3, 50].
@@ III-B1 Preparation Stage
The preparation stage aims to prepare the AI model architecture and data before formal training begins [50]. It includes: (i) model preparation: designing or selecting a base AI model, configuring model size, parameters, architecture, precision settings, and conducting early-phase experimental testing; and (ii) data acquisition and pre-processing: LLMs, which contain billions to hundreds of billions of parameters, require massive training datasets (often terabytes in size) that cover a broad range of content and are collected from diverse sources such as Wikipedia [51], CommonCrawl [52], BookCorpus [53], Reddit posts, GitHub codes, and other text datasets [1]. These raw datasets must be cleaned and preprocessed into training-ready formats (e.g., tokenization, normalization, augmentation). The extract-transform-load pipelines often rely on distributed computing clusters that consume significant electricity during large-scale pre-processing [54]. Well-designed AI models and carefully processed data are critical to the model training efficiency. Empirical results in [55] show that data-centric approaches, such as removing redundant samples from datasets or performing effective feature selection, can drastically reduce the energy consumption of model training by more than 90%, underscoring the importance of developing Green AI [55]. Nevertheless, the preparation stage is flexible and widely dispersed, with the workloads varying substantially depending on factors such as the choice of data processing libraries and tools [56], the deployment environment, and the workflow structure, making its electricity consumption difficult to quantify precisely.
@@ III-B2 Training Stage
The training stage is the most electricity-intensive phase of AI development, particularly for LLMs. During this process, computing hardware operates at near-peak capacity for extended periods, which often runs continuously for days or weeks, resulting in a significant and sustained power draw. For example, training GPT-3 is estimated to have consumed 1.29 GWh of electricity [5]; this intensive process involved approximately 14.8 days of continuous computation on 10,000 NVIDIA V100 GPUs in a Microsoft data center [5]. Furthermore, training GPT-4 required an estimated over 50 GWh of electricity, approximately 40 times more than GPT-3, and equivalent to nearly 0.02% of California’s annual electricity consumption [^1] OpenAI does not officially disclose the energy consumption data for training its GPT models. The figures cited are estimates based on third-party analysis and research. [7, 6]. Similar LLMs, such as Meta’s LLaMA series and Google’s Gemini, require tens of thousands of high-performance GPUs operating in parallel for training, resulting in exceptionally high electricity consumption, often ranging from hundreds to thousands of MWh.
@@
Typical electricity load patterns during the training stage include: (i) Initial load ramp-up phase: at the beginning of training, resource utilization gradually increases as datasets are loaded into memory and distributed across computing nodes, creating a short period of escalating demand [17]. (ii) Constant high demand: computing hardware operates at near-maximum utilization for prolonged periods with very high power load throughout the training stage [20]. (iii) Rapid and large power fluctuations: the alternation between a power-intensive computation phase and a less power-demanding communication phase in the training stage often leads to large power swings [18]. Besides, tasks such as checkpointing, intermediate result saving, and large-scale data transfers between nodes can cause brief but pronounced spikes in electricity consumption; additional transient surges may occur when training is paused and resumed [17, 57]. (iv) High cooling load: sustained full-power operation of GPUs and supporting equipment produces substantial heat, requiring cooling systems to operate at elevated capacity, which significantly increases total energy use. Electricity consumption during the training stage is affected by many factors, such as AI model architecture and training configuration choices, such as batch size and numerical precision requirements. Several strategies have been proposed to improve training efficiency. For example, reference [58] showed that power-capping, which limits the maximum power a GPU can consume, enables a 15% decrease in energy usage with a marginal increase in overall computation time. Other approaches include dynamically adjusting training parameters based on convergence progress and implementing early stopping when performance improvements plateau.
@@ III-B3 Fine-tuning Stage
The fine-tuning stage adapts a pre-trained AI model to a specific downstream task [59], such as legal text analysis, medical report summarization, or customer service chatbots. Since the AI model has already undergone extensive pre-training, fine-tuning typically requires substantially less computation and electricity than training from scratch. Empirical measurements in [60] show that pre-training BERT consumes electricity equivalent to approximately 400 to 45,000 fine-tuning runs, depending on the dataset size and task complexity. Nevertheless, as fine-tuning often involves smaller datasets, frequent evaluation cycles, and hyperparameter tuning, power use tends to be more variable. Rather than maintaining a sustained high baseline, fine-tuning workloads typically exhibit intermittent bursts of GPU utilization followed by lower activity periods.
Typical electricity use patterns during fine-tuning include: (i) moderate average demand, generally lower than full pre-training but varying by model size and dataset; (ii) bursty workloads, with short periods of high consumption during forward/backward passes, validation, and hyperparameter exploration; and (iii) cooling demand variability, with cooling loads fluctuating in step with computational bursts. Although a single fine-tuning run consumes far less energy than pre-training, its much higher frequency across numerous organizations implies that fine-tuning can cumulatively represent a non-negligible share of total lifecycle energy consumption, underscoring the need to enhance its efficiency. For example, reference [61] introduced GreenTrainer, an adaptive fine-tuning algorithm that strategically selects which model parameters to update. This method reduces fine-tuning workload by up to 64% without significant accuracy loss, highlighting the potential to make fine-tuning more energy-efficient [61]. Reference  [62] analyzes the trade-offs between energy consumption and model performance, showing that fine-tuning smaller models such as T5-base, BART-base, or LLaMA3-8B can achieve competitive results while significantly reducing energy usage.
@@ III-B4 Inference Stage
Inference is the process of executing a well-trained AI model to generate outputs in response to user inputs. On a per-query basis, it is the least electricity-intensive stage of the AI model lifecycle. Nevertheless, as inference is performed continuously at scale, often serving millions or billions of requests, its cumulative electricity consumption can exceed that of training. According to [63], compared to 40% for training, inference represents about 60% of total AI energy usage at Google, owing to the billion-user services that rely on AI. Recent estimates suggest inference can account for up to 90% of a model’s total lifecycle energy use [26]. The electricity consumption for a single AI inference query varies widely depending on the model size and architecture, prompt length, and task complexity [26]. Compared to a traditional Google search, which consumes about 0.3 Wh [64], a recent study [26] estimates that GPT-o3 consumes 39.2 Wh, DeepSeek-R1 33.6 Wh, and GPT-4.5 30.5 Wh for processing a long prompt query, whereas GPT-4.1 Nano requires only 0.45 Wh. Reference [65] compares inference energy consumption across a wide range of tasks and shows that image generation is among the most energy-intensive, whereas simpler tasks, such as text classification, consume significantly less electricity.
@@
Unlike the sustained, high load profile of model training, inference workloads are characterized by significant variability. This is driven by fluctuating user requests, diverse query complexities, and the unpredictable nature of real-time interactions, resulting in short, intense bursts of computing load. As a result, typical electricity use patterns in the inference stage include: (i) low average demand per query but high total demand over time due to continuous operation; (ii) bursty and unpredictable spikes from user activity and varying task complexity; and (iii) strong diurnal patterns, with peak usage during business or social activity hours. Supporting such dynamic loads necessitates scalable and flexible IT and power infrastructure capable of rapidly adjusting resources to provide fast response and maintain operational reliability. Energy efficiency can be improved at both the hardware and software levels. Specialized low-power accelerators such as FPGAs and ASICs [66] can reduce energy consumption, while techniques like model quantization, pruning, and knowledge distillation further enhance performance [67, 68]. Emerging architectures such as retrieval augmented generation (RAG) [69] introduce additional complexity by combining model inference with large-scale data retrieval processes. Despite the exponential growth in model parameters, the associated energy cost of inference has been shown to grow sub-linearly, suggesting that efficiency gains can partially offset the demands of larger architectures [70].
@@
Figure 3: Illustration of the patterns of AI computing load (represented by GPU power) during training, fine-tuning, and inference stages. (The data are derived from [16]; refer to Figures 9, 13, 15 and Table III in [16] for detailed data and experimental settings.)
@@ III-C Summary and Discussion
AI computing load is highly energy-intensive, with each stage of the workflow contributing differently to the total electricity consumption. According to [14], the inference stage represents a dominant share of approximately 60% of the energy footprint, while training accounts for 30%, and model preparation and fine-tuning make up the remaining 10%. Moreover, these stages exhibit distinct load patterns. Figure 3 illustrates the typical time-varying electric computing load (measured by GPU power) patterns across different stages of the AI workflow, excluding the data preparation stage due to its typically dispersed energy usage. The electricity load estimates for each stage are derived from [16], where the tasks were executed using GPT-2. Additional experimental details are available in [16]. As shown in Figure 3, all stages exhibit an initial rapid ramp-up phase, during which power consumption rises sharply from a low (near-zero) value to a high operational level. In the training stage, power remains consistently high, characterized by intermittent large rapid load fluctuations, along with persistent high-frequency, small-scale variations, the magnitude of which depends on the underlying hardware and training algorithms [16]. The fine-tuning stage begins with pronounced high-frequency, large-magnitude fluctuations, which gradually decrease in frequency over time; however, moderate fluctuations continue throughout the process. In contrast, the inference stage operates over much shorter durations and exhibits highly variable short bursts of power demand, driven by the stochastic nature of real-time inference requests. This diversity in load profiles underscores the critical need for advanced AI data center designs, grid integration strategies, and power-aware workload scheduling approaches that balance performance, sustainability, and system reliability.
@@ IV-A Long-Term Planning and Interconnection
Over the long-term timescales (years to decades), AI data centers are projected to drive a sustained increase in electricity demand, requiring substantial investment in generation capacity, transmission infrastructure, and distribution equipment to maintain an adequate and reliable electricity supply. According to the IEA report [8], global electricity consumption by data centers is projected to more than double by 2030, reaching approximately 945 TWh, with AI workloads expected to be the primary driver of this growth. However, many regional grids are incapable of accommodating large-scale data centers without extensive transmission and distribution upgrades, which often require 5-10 years for planning, permitting, and construction [71, 22]. Reference [22] assessed the 5‑year projections of AI data-center demand across several bulk power grids and found that resource‑adequacy constraints could severely limit their planned growth, particularly in high‑density clusters such as EirGrid in Ireland and Dominion in the U.S. Moreover, AI data centers tend to concentrate in a limited number of regions, driven by low energy prices, available land and water resources, and supportive policy environments. For example, in the U.S., fifteen states (notably Virginia, Texas, and California) accounted for an estimated 80% of the national data center load in 2023 [14], as shown in Figure 4. This geographical concentration effect further intensifies regional grid stress.
@@
Figure 4: The fifteen U.S. states with the highest data center loads in 2023 [14] (darker colors indicate higher loads).
@@
To address this challenge, coordinated co-planning of AI data centers, the power grid, and generations can reduce costly infrastructure investments, alleviate transmission congestion, enhance operational reliability, and support decarbonization goals. This approach integrates transmission and distribution expansion planning with the strategic siting of AI data centers, as well as the optimal siting and sizing of additional generation and energy storage resources. A number of studies [72, 73, 74, 75] focus on the modeling and optimization of joint planning for conventional data centers and the power grid, while solutions tailored to AI data centers remain largely underexplored. Another key strategy is the co‑location with dedicated local generation, such as renewable generators and nuclear power plants. This scheme can directly supply large stable power sources to high-demand AI workloads, reducing reliance on long-distance electricity transmission and mitigating grid-integration challenges. For example, Amazon Web Services (AWS) has arranged to source approximately 960 MW directly from a nuclear power plant in Pennsylvania to power its data center under a dedicated power purchase agreement [76]. Google and Kairos Power have signed a Master Plant Development Agreement to deploy a fleet of advanced small modular reactors to provide up to 500 MW of clean electricity to Google’s AI data centers [77].
@@
In addition, the grid interconnection policies and regulatory frameworks remain among the most significant bottlenecks to AI data center deployment. For example, in 2025, Texas passed Senate Bill 6 (SB-6) [78], which redefines the interconnection process for large electrical loads, such as data centers exceeding 75 MW, within the ERCOT grid. Under this bill, large load customers must submit interconnection requests through the interconnecting utility and comply with new requirements, including disclosure of similar requests, reporting of on-site backup generation, payment of a minimum \$100k transmission screening study fee, proof of site control, and a financial commitment to cover transmission infrastructure costs. Complex regulatory and permitting workflows, thorough grid integration studies, overloaded interconnection queues, and the need for major grid infrastructure upgrades are causing multiyear delays in data center deployments across regions such as Northern Virginia [79]. Refining interconnection processes and deploying AI and automation tools to accelerate interconnection studies could significantly reduce these delays. Moreover, tariff structure varies across jurisdictions: many data centers fall under smaller regional utilities and are subject to distribution-level tariffs, while hyperscale facilities connect directly to bulk transmission systems, exposing them to wholesale market price signals. As a result, the tariff design and standardization are critical to ensuring cost predictability and sustaining the deployment of AI data centers [80, 81].

### v4

@@ III-B AI Computing Load Patterns at Different Stages
Since AI computing constitutes the dominant share of electricity consumption in AI data centers, this subsection analyzes its load patterns and characteristics across the four main stages of AI model workflows: preparation, training, fine-tuning, and inference [touvron2023llama, liu2024deepseek].
@@ III-B1 Preparation Stage
The preparation stage aims to prepare the AI model architecture and data before formal training begins [liu2024deepseek]. It includes: (i) model preparation: designing or selecting a base AI model, configuring model size, parameters, architecture, precision settings, and conducting early-phase experimental testing; and (ii) data acquisition and pre-processing: LLMs, which contain billions to hundreds of billions of parameters, require massive training datasets (often terabytes in size) that cover a broad range of content and are collected from diverse sources such as Wikipedia [wikipedia], CommonCrawl [commoncrawl], BookCorpus [zhu2015aligning], Reddit posts, GitHub codes, and other text datasets [zhao2023survey]. These raw datasets must be cleaned and preprocessed into training-ready formats (e.g., tokenization, normalization, augmentation). The extract-transform-load pipelines often rely on distributed computing clusters that consume significant electricity during large-scale pre-processing [ai_engineering_wiki]. Well-designed AI models and carefully processed data are critical to the model training efficiency. Empirical results in [verdecchia2022data] show that data-centric approaches, such as removing redundant samples from datasets or performing effective feature selection, can drastically reduce the energy consumption of model training by more than 90%, underscoring the importance of developing Green AI [verdecchia2022data]. Nevertheless, the preparation stage is flexible and widely dispersed, with the workloads varying substantially depending on factors such as the choice of data processing libraries and tools [shanbhag2022energy], the deployment environment, and the workflow structure, making its electricity consumption difficult to quantify precisely.
@@ III-B2 Training Stage
The training stage is the most electricity-intensive phase of AI development, particularly for LLMs. During this process, computing hardware operates at near-peak capacity for extended periods, which often runs continuously for days or weeks, resulting in a significant and sustained power draw. For example, training GPT-3 is estimated to have consumed 1.29 GWh of electricity [patterson2021carbon]; this intensive process involved approximately 14.8 days of continuous computation on 10,000 NVIDIA V100 GPUs in a Microsoft data center [patterson2021carbon]. Furthermore, training GPT-4 required an estimated over 50 GWh of electricity, approximately 40 times more than GPT-3, and equivalent to nearly 0.02% of California’s annual electricity consumption [^1] OpenAI does not officially disclose the energy consumption data for training its GPT models. The figures cited are estimates based on third-party analysis and research. [Cohen2024May, extnet_gpt4_energy]. Similar LLMs, such as Meta’s LLaMA series and Google’s Gemini, require tens of thousands of high-performance GPUs operating in parallel for training, resulting in exceptionally high electricity consumption, often ranging from hundreds to thousands of MWh.
@@
Typical electricity load patterns during the training stage include: (i) Initial load ramp-up phase: at the beginning of training, resource utilization gradually increases as datasets are loaded into memory and distributed across computing nodes, creating a short period of escalating demand [mughees2025forecast]. (ii) Constant high demand: computing hardware operates at near-maximum utilization for prolonged periods with very high power load throughout the training stage [Li2025AILoadDynamics]. (iii) Rapid and large power fluctuations: the alternation between a power-intensive computation phase and a less power-demanding communication phase in the training stage often leads to large power swings [choukse2025power]. Besides, tasks such as checkpointing, intermediate result saving, and large-scale data transfers between nodes can cause brief but pronounced spikes in electricity consumption; additional transient surges may occur when training is paused and resumed [mughees2025forecast, Latif2024EmpiricalGPU]. (iv) High cooling load: sustained full-power operation of GPUs and supporting equipment produces substantial heat, requiring cooling systems to operate at elevated capacity, which significantly increases total energy use. Electricity consumption during the training stage is affected by many factors, such as AI model architecture and training configuration choices, such as batch size and numerical precision requirements. Several strategies have been proposed to improve training efficiency. For example, reference [McDonald2022May] showed that power-capping, which limits the maximum power a GPU can consume, enables a 15% decrease in energy usage with a marginal increase in overall computation time. Other approaches include dynamically adjusting training parameters based on convergence progress and implementing early stopping when performance improvements plateau.
@@ III-B3 Fine-tuning Stage
The fine-tuning stage adapts a pre-trained AI model to a specific downstream task [anisuzzaman2025fine], such as legal text analysis, medical report summarization, or customer service chatbots. Since the AI model has already undergone extensive pre-training, fine-tuning typically requires substantially less computation and electricity than training from scratch. Empirical measurements in [Wang2023FineTuningEnergy] show that pre-training BERT consumes electricity equivalent to approximately 400 to 45,000 fine-tuning runs, depending on the dataset size and task complexity. Nevertheless, as fine-tuning often involves smaller datasets, frequent evaluation cycles, and hyperparameter tuning, power use tends to be more variable. Rather than maintaining a sustained high baseline, fine-tuning workloads typically exhibit intermittent bursts of GPU utilization followed by lower activity periods.
Typical electricity use patterns during fine-tuning include: (i) moderate average demand, generally lower than full pre-training but varying by model size and dataset; (ii) bursty workloads, with short periods of high consumption during forward/backward passes, validation, and hyperparameter exploration; and (iii) cooling demand variability, with cooling loads fluctuating in step with computational bursts. Although a single fine-tuning run consumes far less energy than pre-training, its much higher frequency across numerous organizations implies that fine-tuning can cumulatively represent a non-negligible share of total lifecycle energy consumption, underscoring the need to enhance its efficiency. For example, reference [Huang2023GreenTrainer] introduced GreenTrainer, an adaptive fine-tuning algorithm that strategically selects which model parameters to update. This method reduces fine-tuning workload by up to 64% without significant accuracy loss, highlighting the potential to make fine-tuning more energy-efficient [Huang2023GreenTrainer]. Reference  [Rehman2025Jan] analyzes the trade-offs between energy consumption and model performance, showing that fine-tuning smaller models such as T5-base, BART-base, or LLaMA3-8B can achieve competitive results while significantly reducing energy usage.
@@ III-B4 Inference Stage
Inference is the process of executing a well-trained AI model to generate outputs in response to user inputs. On a per-query basis, it is the least electricity-intensive stage of the AI model lifecycle. Nevertheless, as inference is performed continuously at scale, often serving millions or billions of requests, its cumulative electricity consumption can exceed that of training. According to [patterson2022carbon], compared to 40% for training, inference represents about 60% of total AI energy usage at Google, owing to the billion-user services that rely on AI. Recent estimates suggest inference can account for up to 90% of a model’s total lifecycle energy use [jegham2025hungry]. The electricity consumption for a single AI inference query varies widely depending on the model size and architecture, prompt length, and task complexity [jegham2025hungry]. Compared to a traditional Google search, which consumes about 0.3 Wh [kanoppi2024searchenergy], a recent study [jegham2025hungry] estimates that GPT-o3 consumes 39.2 Wh, DeepSeek-R1 33.6 Wh, and GPT-4.5 30.5 Wh for processing a long prompt query, whereas GPT-4.1 Nano requires only 0.45 Wh. Reference [Luccioni2023Nov] compares inference energy consumption across a wide range of tasks and shows that image generation is among the most energy-intensive, whereas simpler tasks, such as text classification, consume significantly less electricity.
@@
Unlike the sustained, high load profile of model training, inference workloads are characterized by significant variability. This is driven by fluctuating user requests, diverse query complexities, and the unpredictable nature of real-time interactions, resulting in short, intense bursts of computing load. As a result, typical electricity use patterns in the inference stage include: (i) low average demand per query but high total demand over time due to continuous operation; (ii) bursty and unpredictable spikes from user activity and varying task complexity; and (iii) strong diurnal patterns, with peak usage during business or social activity hours. Supporting such dynamic loads necessitates scalable and flexible IT and power infrastructure capable of rapidly adjusting resources to provide fast response and maintain operational reliability. Energy efficiency can be improved at both the hardware and software levels. Specialized low-power accelerators such as FPGAs and ASICs [amara2006fpga] can reduce energy consumption, while techniques like model quantization, pruning, and knowledge distillation further enhance performance [Han2015Pruning, Stock2020Distillation]. Emerging architectures such as retrieval augmented generation (RAG) [fan2024survey] introduce additional complexity by combining model inference with large-scale data retrieval processes. Despite the exponential growth in model parameters, the associated energy cost of inference has been shown to grow sub-linearly, suggesting that efficiency gains can partially offset the demands of larger architectures [Desislavov2021Sep].
@@
Figure 3: Illustration of the patterns of AI computing load (represented by GPU power) during training, fine-tuning, and inference stages. (The data are derived from [Li2024Sep]; refer to Figures 9, 13, 15 and Table III in [Li2024Sep] for detailed data and experimental settings.)
@@ III-C Summary and Discussion
AI computing load is highly energy-intensive, with each stage of the workflow contributing differently to the total electricity consumption. According to [aljbour2024powering], the inference stage represents a dominant share of approximately 60% of the energy footprint, while training accounts for 30%, and model preparation and fine-tuning make up the remaining 10%. Moreover, these stages exhibit distinct load patterns. Figure 3 illustrates the typical time-varying electric computing load (measured by GPU power) patterns across different stages of the AI workflow, excluding the data preparation stage due to its typically dispersed energy usage. The electricity load estimates for each stage are derived from [Li2024Sep], where the tasks were executed using GPT-2. Additional experimental details are available in [Li2024Sep]. As shown in Figure 3, all stages exhibit an initial rapid ramp-up phase, during which power consumption rises sharply from a low (near-zero) value to a high operational level. In the training stage, power remains consistently high, characterized by intermittent large rapid load fluctuations, along with persistent high-frequency, small-scale variations, the magnitude of which depends on the underlying hardware and training algorithms [Li2024Sep]. The fine-tuning stage begins with pronounced high-frequency, large-magnitude fluctuations, which gradually decrease in frequency over time; however, moderate fluctuations continue throughout the process. In contrast, the inference stage operates over much shorter durations and exhibits highly variable short bursts of power demand, driven by the stochastic nature of real-time inference requests. This diversity in load profiles underscores the critical need for advanced AI data center designs, grid integration strategies, and power-aware workload scheduling approaches that balance performance, sustainability, and system reliability.
@@ IV-A Long-Term Planning and Interconnection
Over the long-term timescales (years to decades), AI data centers are projected to drive a sustained increase in electricity demand, requiring substantial investment in generation capacity, transmission infrastructure, and distribution equipment to maintain an adequate and reliable electricity supply. According to the IEA report [iea2024energyai], global electricity consumption by data centers is projected to more than double by 2030, reaching approximately 945 TWh, with AI workloads expected to be the primary driver of this growth. However, many regional grids are incapable of accommodating large-scale data centers without extensive transmission and distribution upgrades, which often require 5-10 years for planning, permitting, and construction [wmeng2025, lin2024exploding]. Reference [lin2024exploding] assessed the 5‑year projections of AI data-center demand across several bulk power grids and found that resource‑adequacy constraints could severely limit their planned growth, particularly in high‑density clusters such as EirGrid in Ireland and Dominion in the U.S. Moreover, AI data centers tend to concentrate in a limited number of regions, driven by low energy prices, available land and water resources, and supportive policy environments. For example, in the U.S., fifteen states (notably Virginia, Texas, and California) accounted for an estimated 80% of the national data center load in 2023 [aljbour2024powering], as shown in Figure 4. This geographical concentration effect further intensifies regional grid stress.
@@
Figure 4: The fifteen U.S. states with the highest data center loads in 2023 [aljbour2024powering] (darker colors indicate higher loads).
@@
To address this challenge, coordinated co-planning of AI data centers, the power grid, and generations can reduce costly infrastructure investments, alleviate transmission congestion, enhance operational reliability, and support decarbonization goals. This approach integrates transmission and distribution expansion planning with the strategic siting of AI data centers, as well as the optimal siting and sizing of additional generation and energy storage resources. A number of studies [chen2024joint, guo2021integrated, vafamehr2017framework, zhang2024multi] focus on the modeling and optimization of joint planning for conventional data centers and the power grid, while solutions tailored to AI data centers remain largely underexplored. Another key strategy is the co‑location with dedicated local generation, such as renewable generators and nuclear power plants. This scheme can directly supply large stable power sources to high-demand AI workloads, reducing reliance on long-distance electricity transmission and mitigating grid-integration challenges. For example, Amazon Web Services (AWS) has arranged to source approximately 960 MW directly from a nuclear power plant in Pennsylvania to power its data center under a dedicated power purchase agreement [aws2024]. Google and Kairos Power have signed a Master Plant Development Agreement to deploy a fleet of advanced small modular reactors to provide up to 500 MW of clean electricity to Google’s AI data centers [patel2024_google_kairos].
@@
In addition, the grid interconnection policies and regulatory frameworks remain among the most significant bottlenecks to AI data center deployment. For example, in 2025, Texas passed Senate Bill 6 (SB-6) [TXSB6_2025], which redefines the interconnection process for large electrical loads, such as data centers exceeding 75 MW, within the ERCOT grid. Under this bill, large load customers must submit interconnection requests through the interconnecting utility and comply with new requirements, including disclosure of similar requests, reporting of on-site backup generation, payment of a minimum \$100k transmission screening study fee, proof of site control, and a financial commitment to cover transmission infrastructure costs. Complex regulatory and permitting workflows, thorough grid integration studies, overloaded interconnection queues, and the need for major grid infrastructure upgrades are causing multiyear delays in data center deployments across regions such as Northern Virginia [FTI2024PowerTrends]. Refining interconnection processes and deploying AI and automation tools to accelerate interconnection studies could significantly reduce these delays. Moreover, tariff structure varies across jurisdictions: many data centers fall under smaller regional utilities and are subject to distribution-level tariffs, while hyperscale facilities connect directly to bulk transmission systems, exposing them to wholesale market price signals. As a result, the tariff design and standardization are critical to ensuring cost predictability and sustaining the deployment of AI data centers [satchwell2025electricity, CRA2024TariffDesign].